# A Comparative Study of Conventional and Tripolar EEG for High-Performance Reach-to-Grasp BCI Systems

### v1

@@ Abstract
Objective. This study aims to enhance brain-computer interface (BCI) applications for individuals with motor impairments by comprehensively comparing the effectiveness of noninvasive tripolar EEG (tEEG) with conventional EEG technology. The focus is on accurately interpreting and decoding various grasping movements, such as power grasp and precision grasp. The goal is to determine which EEG technology is more effective in processing and translating grasp-related neural signals. Approach. The approach involved conducting an experiment on ten healthy participants who performed two distinct reach-and-grasp movements: power grasp and precision grasp, with a ’no-movement’ condition serving as the baseline. Our research presents a thorough comparison between EEG and tEEG in decoding grasping movements. This comparison spans several key parameters, including signal-to-noise ratio (SNR), spatial resolution via functional connectivity, event-related potentials (ERPs), and wavelet time-frequency analysis. Additionally, our study involved extracting and analyzing statistical features from the wavelet coefficients, and both binary and multiclass classification methods were employed. Four machine learning algorithms—Random Forest (RF), Support Vector Machine (SVM), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA)—were used to evaluate the decoding accuracies. Main results. Our results indicated that tEEG demonstrated superior performance over conventional EEG in various aspects. This included a higher signal-to-noise ratio, enhanced spatial resolution, and more informative data in ERPs and wavelet time-frequency analysis. The use of tEEG led to notable improvements in decoding accuracy for differentiating movement types. Specifically, tEEG achieved around 90% accuracy in binary and 75.97% for multiclass classification. These results are markedly better than those from standard EEG, which recorded a maximum of 77.85% and 61.27% in similar tasks, respectively. These findings highlight the superior effectiveness of tEEG over conventional EEG in decoding grasp types and its competitive or superior performance in complex classifications compared with existing research.
Keywords: brain-computer-interface, motor impairments, tripolar EEG (tEEG), conventional EEG, grasping movements, machine learning
@@ 1 Introduction
Brain-computer interfaces (BCIs) represent a transformative technology, creating a direct communication link between neural activity and computer systems [1, 2, 3]. They offer substantial benefits to individuals with severe motor impairments, especially those resulting from spinal cord injuries, enhancing their quality of life through control over assistive devices and technologies. A key and challenging aspect in BCI applications is decoding human grasping intentions. This isn’t just about motor function; it integrates sensory perceptions, motor planning, and cognitive intentions [4]. The significance of accurately decoding these intentions extends beyond scientific interest, holding the potential to revolutionize assistive technologies and facilitate natural interactions for BCI users.
@@
The act of reaching and grasping is central to human motor functions and crucial for daily tasks. This involves various brain areas: the parietal regions, including the anterior intraparietal sulcus, the premotor regions like the ventral premotor cortex, and the primary motor cortex [5, 6]. While grasping plays a pivotal role in our interactions, decoding its underlying complex neuronal mechanisms remains a primary objective in neuroscience [7]. Achieving this can transform numerous domains: from enhancing rehabilitation medicine [8, 9, 10, 11] and advancing assistive technology [12, 13, 14] to designing intuitive brain-machine interfaces [15, 16, 17]. Grasping requires precise muscle and joint coordination, adapting to the object’s properties. Unraveling such intricate actions from brain signals is a challenging task. Many studies have explored the brain patterns linked to grasping actions in non-human primates [18, 19, 20] and humans using different acquisition modalities including invasive, such as electrocorticography (ECoG) [21] and noninvasive, such as functional magnetic resonance imaging (fMRI) [22, 23]. Additionally, numerous studies have demonstrated that noninvasive Electroencephalogram (EEG) signals can effectively decode reach-and-grasp actions. In this context, machine learning algorithms are pivotal in decoding complex neural data. Deep learning models like Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) offer advanced analysis capabilities but usually require large datasets, which can be challenging to collect from human participants [24]. Therefore, many studies prefer traditional models such as SVMs, Decision Trees, and LDA. These traditional algorithms are more manageable with smaller datasets and still significantly improve the accuracy of movement decoding and prediction [25]. Iturrate et al [26] utilized EEG to distinguish brain activation patterns during self-paced grasping tasks involving both precision and power grips, achieving 70% accurate single-trial decoding by using shrinkage LDA (sLDA). Schwarz et al. [27] demonstrated that it’s possible to distinguish between three types of reach-and-grasp actions—palmar, pincer, and lateral grasp—using EEG neural correlates. The study achieved classification accuracies of up to 72.4% between grasp types and 93.5% for grasps versus no-movement conditions, supporting the potential use of this technique for more intuitive control of neuroprostheses. In another study by Xu et al [28], five distinct reach-and-grasp movements were decoded using movement-related cortical potentials (MRCPs) from noninvasive EEG signals. The movements included palmar, pinch, push, twist, and plug grasp. The study achieved a peak average accuracy of about 75% for grasping vs. no-movement conditions.
@@
Despite the numerous studies utilizing EEG for decoding grasping movements, EEG often suffers from a low spatial resolution, complicating the precise localization of the source of neural activity due to the blurring effects primarily from different conductivities of the volume conductor [29]. It is also vulnerable to physiological disturbances caused by muscle contractions, eye movements, and heart activity, as well as technical artifacts like power-line noises or fluctuations in electrode impedances. These limitations directly impacts the decoding accuracy of motor intentions. Such a limitation emphasizes the need to explore advanced methods aimed at improving the spatial resolution of EEG data and increasing the signal-to-noise ratio. A promising solution is the use of tEEG, a novel technique involving an additional local reference electrode. The tripolar concentric ring electrode (TCRE) is roughly the same size as a conventional electrode. However, it consists of a central conductive disc with two electrically isolated rings surrounding it. The output signal from the TCRE is derived from a weighted sum of the outer ring subtracted from the central disc, which is then subtracted from sixteen times the difference between the inner ring and the central disc [30]. Compared to conventional disc electrode EEG, TCRE offers superior performance, with roughly 2.5 times higher spatial selectivity, 3.7 times higher signal-to-noise ratio, and approximately 12 times lower mutual information [31, 32]. There are other studies that demonstrated tEEG could effectively locate high-frequency activity in epilepsy patients, predominantly in the seizure onset zone, suggesting its potential utility in identifying epileptic brain regions [33, 34]. Also in [35], authors illustrated that tEEG significantly outperformed conventional disc electrodes in distinguishing individual finger movements, based on movement-related potentials (MRPs). Despite the potential benefits of tEEG, the use of tEEG in the realm of BCIs, particularly for decoding human grasping actions, remains largely unexplored, underlining the novelty and potential value of our study.
In this study, we embark on two pivotal explorations. Firstly, we conduct an in-depth comparison between conventional EEG and tEEG, examining parameters such as SNR, functional connectivity, and ERPs. Our visual illustrations, using topographical maps, aim to underscore the heightened spatial resolution of tEEG with respect to conventional EEG. As a subsequent step, we leverage the precision of wavelets to extract crucial features from the captured brain data. With the assistance of machine learning models, our overall goal is to enhance the decoding accuracy of grasping intentions. Through these endeavors, our research strives to emphasize the superiority of tEEG as a novel non-invasive neuroimaging technique in capturing brain signals and their mutual connectivities as well as the innovative potential of using wavelets technique in interpreting them.
@@ 2 Materials and Methods
A total of ten healthy volunteers, aged between 22 and 35 years consisting of five males and five females, participated in this study. The study was non-invasive and received approval from the Institutional Review Board (IRB). All participants had no known prior medical conditions and possessed normal or corrected-to-normal vision. They were thoroughly briefed about the study’s objectives. Before initiating any recordings, each participant provided informed consent, ensuring they understood the research’s purpose and procedures. The tasks they undertook involved performing grasping movements, specifically targeting objects such as a bottle or pen.
@@ 2.1 Experimental Setup and Paradigm
To thoroughly study the neural activity associated with natural vision-driven grasping tasks, we established a novel and comprehensive experimental setup and study protocol. Our experimental setup consists of a 3D-engineered, motorized turntable divided into three segments for displaying two objects (a bottle and a pen) as well as an empty space for a no-object scenario. A notable advancement in our system compared to other studies is the incorporation of innovative smart eyeglasses equipped with a special film that alternates between clear and opaque. This feature crucially reduces any potential anticipatory bias. Participants were comfortably seated, hands facing down, and placed 30 cm from the object’s center (Figure 1a). As illustrated in Figure 1c, following a three-second observation period accompanied by a buzzing sound signal, the participants commenced the grasping phase. The turntable then rotated to the next object or no-object condition randomly, during which the eyeglasses turned opaque, thus preparing for the subsequent phase.
@@ 2.2 Data Acquisition
EEG data acquisition was realized using the Unicorn Hybrid Black Bluetooth amplifier, capturing data at a rate of 250 samples per second per channel. Prior to initiating each session, skin-to-electrode impedance was assured to be below 10k to optimize data integrity. While Alzahrani’s work [35] demonstrated that only a single TCRE electrode at the C3 motor area location was sufficient for classifying imaginary finger movements, our study extends this approach by also considering additional brain regions such as parietal, central, and frontal areas, which play distinct and vital roles in the various stages of grasping tasks, from planning to execution [36]. To this end, we have strategically placed four TCREs at the P3, C3, C4, and F5 locations based on the 10-20 system, concurrently capturing tEEG and conventional EEG data [37]. Reference and ground electrodes were fixed at the left and right mastoids, respectively. Each participant underwent 50 trials for each condition, allowing us to gather a detailed and extensive dataset for further analysis.
@@ 2.3 Preprocessing
Data analysis was diligently executed utilizing custom Python scripts, leveraging functions from the ’mne’ package [38]. Initially, the data underwent a high-pass filter at 1Hz to minimize signal drifts [39]. Following this foundational step, the recordings were bandpass filtered using a Finite Impulse Response (FIR) filter of order 20, spanning a frequency range of 1Hz—60Hz [40, 41]. This was paired with a notch filter targeting the elimination of the omnipresent 60Hz line noise.
Post the spectral filtering, wavelet denoising, an advanced signal processing technique, was implemented to further refine the quality of the EEG recordings. As shown by the authors in [42], wavelet denoising can be beneficial in EEG signals by removing noise while preserving salient features associated with underlying neuronal activities. For our denoising procedure, we utilized the Daubechies wavelet, specifically the ’db8’ waveform, known for its excellent localization properties in both the time and frequency domains [43]. The ’db8’ wavelet has been identified as the most effective wavelet for noise reduction in EEG signals when analyzing data from healthy subjects [44]. This particular wavelet is distinguished by its superior ability to eliminate extraneous noise, enhancing the clarity and reliability of EEG readings[45].
With the denoised signals in place, the subtraction of the mean reference and a detrending process was performed, aiming to counteract the introduction of spurious trends in the spectral data [46]. To ensure a consistent scale across data, the signals were subjected to Z-score normalization. This technique is essential for ensuring that EEG data is uniform across different measurements, which significantly contributes to improving the accuracy and effectiveness of both traditional machine learning and advanced deep learning models [47].
For a focused examination of movement-related activity, the data were meticulously segmented to capture the temporal window spanning 1 second before to 1 second after movement onset. Culminating the preprocessing pipeline, an Independent Component Analysis (ICA) was administered [48, 49]. This sophisticated technique identified and eliminated components tainted by physiological or non-physiological noise extraneous to EEG signals, ensuring the highest fidelity in data quality for subsequent analytical endeavors.
@@ 2.4 Comparative Analysis
The foundational aspect of this study revolves around a comprehensive comparison between conventional EEG and tEEG technologies. Before delving into classification-based analyses, it is crucial to discern the inherent differences and potential advantages one technology may offer over the other. This section presents the methodologies employed to undertake such comparative analyses, focusing on four pivotal metrics: SNR, functional connectivity, time domain analysis using Event-Related Potentials (ERP), and Time-Frequency analysis using wavelets.
2.4.1 Signal-to-Noise Ratio (SNR)
In evaluating the efficacy of our preprocessing steps, we analyzed the SNR for both EEG and tEEG configurations across multiple channels. This analysis aimed to verify if the tEEG setup indeed offers enhanced signal quality compared to the standard EEG. For calculating SNR, we compared the wavelet-denoised signal to the noise, which is determined as the difference between the original signal and its denoised version. The SNR is computed using the formula:
Here, RMS represents the Root Mean Square. In the context of a continuous waveform, the RMS is the square root of the mean power of the signal. For a discrete sequence, it refers to the square root of the sum of the squares over the total data points. Both EEG and tEEG data underwent SNR computation across all channels. It’s noteworthy to mention that the presented results represent grand averages accumulated from all subjects, ensuring a comprehensive overview. The data used for this analysis was specifically segmented around the peri-movement time frame.
2.4.2 Frequency-Domain Functional Connectivity Analysis using Coherence
@@
2.4.3 Event-Related Potentials (ERP) Analysis
Event-related potentials (ERPs) are voltage fluctuations resulting from neural activations, closely tied to specific sensory, cognitive, or motor events. They offer a non-invasive method to study the processing of external and internal stimuli in the brain. This section seeks to compare the ERP outcomes from conventional EEG and the emerging tEEG during movement onset. For this specific comparison, ERPs were analyzed in the context of movement onset. We focused on three primary conditions: No Movement, Power Grasp, and Precision Grasp. By aligning the EEG data with the exact moment of movement initiation, we sought to ascertain the brain’s immediate response to these motor activities.
2.4.4 Time-Frequency Analysis Using Wavelets
Building upon the foundational analyses of SNR, ERPs, and functional connectivity, we extend our methodological framework to encompass time-frequency analysis using wavelets. While features derived from the time domain, such as ERPs, provide valuable insights into neural dynamics, they offer a limited perspective on the intricate oscillatory nature of brain signals. To capture a more comprehensive feature map, it is imperative to consider both the time and frequency characteristics of the EEG and tEEG data.
@@ 2.4.4 Time-Frequency Analysis Using Wavelets
In our computational approach, we first determined the power of the wavelet coefficients by taking their absolute values and squaring them. This step was crucial for emphasizing the energy content at each frequency and time point. Subsequently, we averaged these power values across all trials to yield a mean spectral density for each channel. The resulting average power spectra were visualized as heatmaps, which allowed for a direct comparison between the EEG and tEEG data. These visual representations were scaled to facilitate a clear comparison, with the color intensity indicating the power at each time-frequency point.
The average power spectra, visualized as heatmaps, serve not only as a tool for feature extraction but also as a means for a detailed comparison between EEG and tEEG. The heatmaps facilitate a side-by-side evaluation of how EEG and tEEG represent neural dynamics across different frequency bands and time intervals. This comparative analysis is crucial, as it allows us to discern which modality provides a richer and more detailed depiction of brain activity, thereby offering superior features for the classification of motor tasks.

### v2

@@ Abstract
Objective: This study aims to enhance brain-computer interface (BCI) applications for individuals with motor impairments by comparing the effectiveness of noninvasive tripolar concentric ring electrode electroencephalography (tEEG) with conventional electroencephalography (EEG) technology. The goal is to determine which EEG technology is more effective in measuring and decoding different grasp-related neural signals. Methods: The approach involves experimenting on ten healthy participants who performed two distinct reach-and-grasp movements: power grasp and precision grasp, with a no-movement condition serving as the baseline. Our research compares EEG and tEEG in decoding grasping movements, focusing on signal-to-noise ratio (SNR), spatial resolution, and wavelet time-frequency analysis. Additionally, our study involved extracting and analyzing statistical features from the wavelet coefficients, and both binary and multiclass classification methods were employed. Four machine learning algorithms—Random Forest (RF), Support Vector Machine (SVM), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA)—were used to evaluate the decoding accuracies. Results: Our results indicated that compared with conventional EEG, tEEG yielded higher SNR, finer spatial resolution, and stronger wavelet power spectra. These advantages translated to superior decoding accuracies: tEEG reached around 90.00% accuracy in binary classification and 75.97% in multiclass tasks, versus 77.85% and 61.27% for conventional EEG. Conclusion: tEEG provides richer and cleaner neural information than conventional EEG, enabling significantly better differentiation of grasp types. Significance: These findings position tEEG as a promising alternative to conventional EEG for next-generation BCIs aimed at restoring or augmenting motor function in people with upper limb movement disabilities.
Brain-computer-interface, conventional EEG, grasping movements, machine learning, motor impairments, tripolar EEG.
@@ 1 Introduction
The development of brain-computer interfaces (BCIs) introduces a groundbreaking technology that facilitates direct interaction between the human brain and computer systems [schalk2004bci2000, abiri2019comprehensive]. They offer substantial benefits to individuals with severe motor impairments, especially those resulting from spinal cord injuries, enhancing their quality of life through control over assistive devices and technologies [rabiee2024streams, ghafoori2025novel]. A key and challenging aspect in BCI applications is decoding grasp-related neural activity. This isn’t just about motor function; it integrates sensory perceptions, motor planning, and cognitive intentions [castiello2008cortical]. The significance of accurately decoding these activities extends beyond scientific interest, holding the potential to revolutionize assistive technologies and facilitate natural interactions for BCI users.
@@
The act of reaching and grasping is central to human motor functions and crucial for daily tasks. This involves various brain areas: the parietal regions, including the anterior intraparietal sulcus, the premotor regions like the ventral premotor cortex, and the primary motor cortex [grafton2010cognitive, gallivan2013decoding]. While grasping plays a pivotal role in our interactions, decoding its underlying complex neuronal mechanisms remains a primary objective in neuroscience [castiello2005neuroscience]. Achieving this can transform numerous domains: from enhancing rehabilitation medicine [park2013rehabilitation, lambercy2009rehabilitation, thrasher2008rehabilitation] and advancing assistive technology [grimm2016closed, rupp2014brain, millan2010combining, rabiee2024wavelet, ghafoori2024bispectrum] to designing intuitive brain-machine interfaces [hazrati2010online, chang2019architecture]. Grasping requires precise muscle and joint coordination, adapting to the object’s properties. Unraveling such intricate actions from brain signals is a challenging task. Many studies have explored the brain patterns linked to grasping actions in non-human primates [donoghue1998neural, schaffelhofer2016object, michaels2018population] and humans using different acquisition modalities including invasive, such as electrocorticography (ECoG) [jiang2020power] and noninvasive, such as functional magnetic resonance imaging (fMRI) [fabbri2016disentangling, perini2020neural]. Additionally, numerous studies have demonstrated that noninvasive Electroencephalogram (EEG) signals can effectively decode reach-and-grasp actions. In this context, machine learning algorithms are pivotal in decoding complex neural data. Deep learning models like Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) offer advanced analysis capabilities but usually require large datasets, which can be challenging to collect from human participants [alzubaidi2023survey]. Therefore, many studies prefer traditional models such as SVMs, Decision Trees, and LDA. These traditional algorithms are more manageable with smaller datasets and still significantly improve the accuracy of movement decoding and prediction [saeidi2021neural]. Iturrate et al [iturrate2018human] utilized EEG to distinguish brain activation patterns during self-paced grasping tasks involving both precision and power grips, achieving 70% accurate single-trial decoding by using shrinkage LDA (sLDA). Schwarz et al. [schwarz2017decoding] demonstrated that it’s possible to distinguish between three types of reach-and-grasp actions—palmar, pincer, and lateral grasp—using EEG neural correlates. The study achieved classification accuracies of up to 72.4% between grasp types and 93.5% for grasps versus no-movement conditions, supporting the potential use of this technique for more intuitive control of neuroprostheses. In another study by Xu et al [xu2021decoding], five distinct reach-and-grasp movements were decoded using movement-related cortical potentials (MRCPs) from noninvasive EEG signals. The movements included palmar, pinch, push, twist, and plug grasp. The study achieved a peak average accuracy of about 75% for grasping vs. no-movement conditions.
@@
Despite the numerous studies utilizing EEG for decoding grasping movements, EEG often suffers from a low spatial resolution, complicating the precise localization of the source of neural activity due to the blurring effects primarily from different conductivities of the volume conductor [nunez1994theoretical]. It is also vulnerable to physiological disturbances caused by muscle contractions, eye movements, and heart activity, as well as technical artifacts like power-line noises or fluctuations in electrode impedances [jiang2019removal]. These limitations directly impact the decoding of grasp-related neural activity. Such a limitation emphasizes the need to explore advanced methods aimed at improving the spatial resolution of EEG data and increasing the signal-to-noise ratio. A promising solution is the use of tEEG, a novel technique involving an additional local reference electrode. The tripolar concentric ring electrode (TCRE) is roughly the same size as a conventional electrode. However, it consists of a central conductive disc with two electrically isolated rings surrounding it. The output signal from the TCRE is derived from a weighted sum of the outer ring subtracted from the central disc, which is then subtracted from sixteen times the difference between the inner ring and the central disc [besio2006tri]. Compared to conventional disc electrode EEG, TCRE offers superior performance, with roughly 2.5 times higher spatial selectivity, 3.7 times higher signal-to-noise ratio, and approximately 12 times lower mutual information [koka2007improvement, aghaei2020tripolar]. There are other studies [toole2019source, makeyev2017recent] that demonstrated tEEG could effectively locate high-frequency activity in epilepsy patients, predominantly in the seizure onset zone, suggesting its potential utility in identifying epileptic brain regions. Also in [alzahrani2021comparison], authors illustrated that tEEG significantly outperformed conventional disc electrodes in distinguishing individual finger movements, based on movement-related potentials (MRPs). Despite the potential benefits of tEEG, the use of tEEG in the realm of BCIs, particularly for decoding human grasping actions, remains largely unexplored.
In this study, we address critical challenges in decoding grasp-related neural activity and propose novel approaches to improve brain-computer interface (BCI) performance. The main contributions of this work are summarized as follows:
This research systematically compares tEEG and conventional EEG technologies for grasp-related tasks, demonstrating significant improvements not only in SNR and spatial resolution but also in decoding accuracy with tEEG, particularly in both binary and multiclass classification scenarios.
The use of wavelet-based techniques alongside tEEG electrodes improves the interpretation of neural activity and decoding accuracy, offering deeper insights into the dynamic nature of grasp-related brain signals.
Despite utilizing only four electrodes, traditional machine learning algorithms, and a cost-effective EEG headset, the study achieves higher classification accuracies compared to existing methods, demonstrating the feasibility of achieving high performance with minimal hardware and computational resources.
The features extracted using wavelet-based techniques and tEEG signals are demonstrated to be rich and robust, enabling consistent performance across various machine learning algorithms.
@@ 2 Materials and Methods
A total of ten healthy volunteers, aged between 22 and 35 years consisting of five males and five females, participated in this study. To determine the requisite sample size for our study, a preliminary power analysis was conducted using G*Power [kang2021sample], based on our pilot data. This analysis aimed to identify the number of participants needed to detect a statistically significant difference in decoding accuracies between two conditions, and the results indicated that a total of 10 participants would be sufficient for this purpose. The study was non-invasive and received approval from the University of Rhode Island Institutional Review Board (IRB, #1944644-4 on January 4, 2024). All participants had no known prior medical conditions and possessed normal or corrected-to-normal vision. They were thoroughly briefed about the study’s objectives. Before initiating any recordings, each participant provided informed consent, ensuring they understood the research’s purpose and procedures.
@@ 2.1 Experimental Setup and Paradigm
To study the neural activity associated with natural vision-driven grasping tasks, we established a novel and comprehensive experimental setup and study protocol. As illustrated in Figure 2A, the setup includes tripolar electrodes, a t-interface pre-amplifier, a g.tec Unicorn Hybrid Black amplifier, an Arduino, and custom software developed to manage the experiment and integrate various hardware components. A key feature of the setup is a 3D-engineered, motorized turntable divided into three segments, which display two objects (a bottle and a pen) or an empty space for a no-object scenario. Participants were seated comfortably with their hands facing down, positioned 30 cm from the object’s center. Following a three-second observation period accompanied by a buzzing sound signal, participants initiated the grasping phase. The turntable then rotated randomly to the next condition—object A, object B, or no-object—while their eyeglasses turned opaque to prepare for the subsequent phase (Figure 2C, 2D).
A notable advancement in our experimental design is the incorporation of these innovative smart eyeglasses, equipped with a special film that alternates between clear and opaque states developed in our previous study [cetera2024classification]. This feature is specifically designed to reduce potential anticipatory bias by obscuring the participants’ view of the objects until the experimental condition is fully set. This addition addresses a known challenge in studies of neural activity, as prior findings, such as those discussing the Fehrer-Raab effect [fehrer1962reaction, breitmeyer2006visual, schurger2010reproducibility], suggest that neural activities can begin before a target becomes fully visible, influenced by prior context or subtle visual cues. In our setup, neural activities might initiate even before the turntable completes its rotation, as participants could anticipate the upcoming condition. By leveraging visual masking through the eyeglasses, our approach minimizes this anticipatory activity, ensuring that the recorded neural signals more accurately reflect task-relevant processes associated with the actual execution phase.
@@ 2.2 Data Acquisition
EEG data acquisition was realized using the Unicorn Hybrid Black Bluetooth amplifier, capturing data at a rate of 250 samples per second per channel. Prior to initiating each session, skin-to-electrode impedance was assured to be below 10k by impedance meter to optimize data integrity. While Alzahrani’s work [alzahrani2021comparison] demonstrated that only a single TCRE electrode at the C3 motor area location was sufficient for classifying imaginary finger movements, our study extends this approach by also considering additional brain regions such as parietal, central, and frontal areas, which play distinct and vital roles in the various stages of grasping tasks, from planning to execution [vingerhoets2014contribution]. To this end, we have placed four TCREs at the P3, C3, C4, and F5 locations based on the 10-20 system as illustrated in Figure 2B, concurrently capturing tEEG and conventional EEG data [makeyev2013emulating]. Reference and ground electrodes were fixed at the left and right mastoids, respectively. Each participant underwent 50 trials for each condition, allowing us to gather a detailed and extensive dataset for further analysis.
@@ 2.3 Preprocessing
Data analysis was executed utilizing custom Python scripts, leveraging functions from the mne package [gramfort2013meg]. Initially, the data were bandpass filtered using a Finite Impulse Response (FIR) filter of order 20, spanning a frequency range of 1Hz—60Hz [widmann2015digital, widmann2012filter]. Post the spectral filtering, wavelet denoising, an advanced signal processing technique, was implemented to further refine the quality of the EEG recordings. As shown by the authors in [carmona1994wavelet], wavelet denoising can be beneficial in EEG signals by removing noise while preserving salient features associated with underlying neuronal activities. For our denoising procedure, we utilized the Daubechies wavelet, specifically the db8 waveform, known for its localization properties in both the time and frequency domains [daubechies1992ten]. The db8 wavelet has been identified as the most effective wavelet for noise reduction in EEG signals when analyzing data from healthy subjects [mamun2013effectiveness]. This particular wavelet is distinguished by its ability to eliminate extraneous noise, enhancing the clarity and reliability of EEG readings[asaduzzaman2010study].
With the denoised signals in place, the subtraction of the mean reference and a detrending process was performed, aiming to counteract the introduction of spurious trends in the spectral data [de2018robust]. To ensure a consistent scale across data, the signals were subjected to Z-score normalization. This technique is essential for ensuring that EEG data is uniform across different measurements, which significantly contributes to improving the accuracy and effectiveness of both traditional machine learning and advanced deep learning models [shoeibi2021automatic].
For a focused examination of movement-related activity, the data were meticulously segmented to capture the temporal window spanning 1 second before to 1 second after movement onset. Culminating the preprocessing pipeline, an Independent Component Analysis (ICA) was administered [hyvarinen2000independent, bell1995information]. This technique identified and eliminated components tainted by physiological or non-physiological noise extraneous to EEG signals, for subsequent data analysis.
@@ 2.4 Comparative Analysis
The foundational aspect of this study revolves around a comprehensive comparison between conventional EEG and tEEG technologies. Before delving into classification-based analyses, it is crucial to discern the inherent differences and potential advantages one technology may offer over the other. This section presents the methodologies employed to undertake such comparative analyses, focusing on three pivotal techniques: SNR, functional connectivity, and Time-Frequency analysis using wavelets.
Signal-to-Noise Ratio (SNR)
In evaluating the efficacy of our preprocessing steps, we analyzed the SNR for both EEG and tEEG configurations across multiple channels while participants performed the grasping tasks. This analysis aimed to verify if the tEEG setup indeed offers enhanced signal quality compared to the standard EEG. For calculating SNR, we compared the wavelet-denoised signal to the noise, which is determined as the difference between the original signal and its denoised version. The SNR is computed using the formula:
Here, RMS represents the Root Mean Square. In the context of a continuous waveform, the RMS is the square root of the mean power of the signal. For a discrete sequence, it refers to the square root of the sum of the squares over the total data points. Both EEG and tEEG data underwent SNR computation across all channels. It is noteworthy to mention that the presented results represent grand averages accumulated from all subjects, ensuring a comprehensive overview. The data used for this analysis was specifically segmented around the peri-movement time frame.
Frequency-Domain Functional Connectivity Analysis
@@
Time-Frequency Analysis
Building upon the foundational analyses of SNR and functional connectivity, we extend our methodological framework to encompass time-frequency analysis by using wavelets. While features derived from the time domain, provide valuable insights into neural dynamics, they offer a limited perspective on the intricate oscillatory nature of brain signals [cohen2014analyzing]. To capture a more comprehensive feature map, it is imperative to consider both the time and frequency characteristics of the EEG and tEEG data.
@@
In our computational approach, we first determined the power of the wavelet coefficients by calculating the absolute value to find the energy content at each frequency and time point. Subsequently, we averaged these power values across all trials to yield a mean spectral density for each channel. The resulting average power spectra were visualized as heatmaps, which allowed for a direct comparison between the EEG and tEEG data. These visual representations were scaled to facilitate a clear comparison, with the color intensity indicating the power at each time-frequency point.
The average power spectra, visualized as heatmaps, serve not only as a tool for feature extraction but also as a means for a detailed comparison between EEG and tEEG. The heatmaps facilitate a side-by-side evaluation of how EEG and tEEG represent neural dynamics across different frequency bands and time intervals. This comparative analysis allows us to discern which modality provides a richer and more detailed depiction of brain activity, thereby offering more informative features for the classification of motor tasks.